OpenStudy (anonymous):

does g(x) =(-x)^2 reflect in the x axis?

OpenStudy (anonymous):

(-x)^2 is just x^2. So what is your question?

OpenStudy (anonymous):

I am trying to figure out what its doing on a graph

OpenStudy (anonymous):

I think it reflects in the x axis?

OpenStudy (anonymous):

x^2 or y = x^2 is a parabola..U-shaped with a vertex at the origin (0,0).I still don't understand what you mean by "what its doing on a graph"?

OpenStudy (anonymous):

If you reflect y = x^2 in the x-axis, then it will be an upside down u shaped parabola and its equation is y = -x^2. But nearly all curves can be reflected in the x-axis. Am I missing what's bothering you?

OpenStudy (anonymous):

well here is what it says I have to do: Matching. The left-hand column contains equations that represent transformations of f(x) = x^2. Match the equations on the left with the description on the right of how to obtain the graph of g from the graph of f.

OpenStudy (anonymous):

then there are a bunch of options and I get all the other ones now except for this. unless I am right because I think the parabola reflects in the x acis?

OpenStudy (anonymous):

As I said, if you reflect the parabola in the x-axis you will also get a parabola whose equation is y = -x^2. I will try drawing a diagram.

OpenStudy (anonymous):

oh it reflects in the y acis!

OpenStudy (anonymous):

axis

OpenStudy (anonymous):



OpenStudy (anonymous):

because: fx− Reflect ()fx in the y-axis Change every x in the function to -x

OpenStudy (anonymous):

I just showed you y = x^2 and its reflection in the x-axis, called y = -x^2

OpenStudy (anonymous):

that's none of my options for answers

OpenStudy (anonymous):

but it would reflect in the Y axis then right? because that's what it looks like.

OpenStudy (ranga):

If f(x) = x^2 and you get g(x) by replacing x with -x, that is, g(x) = f(-x) then g(x) is a reflection over the y axis. In this case the curves f(x) and g(x) will look the same.